vertical angle of an isosceles triangle is25degress 17minutes.express the base angle in circular measure

Answer by cleomenius(959)

First, we must find the base angle, as we are given the vertex angle
180 - 25 degree 17 minutes / 2 = 77.215 degrees.
The question is asking for the answer in circular measure, or radions
77.215 * pi/180 = 0.429 * pi, or approx. 1.348 radians.
Cleomenius.